Free cafoa valuation) The Brown Corporation is viewed as a posible takeover turget y Cicron, Inc. Curentdy, Brown uses 20 percent debt in its capital structure, but Cicron plans to increase the debt ratio to 25 percent if the acquisition is consummated. The after-tax cost of debt apital for Brown is estimated to be 8 percent,which holds constant under either capital structure. The cost of equity fer the acquisition is eapected to be 22 percent. The current mar- ket value of Browns outstanding debt is $75 milin, all of which will be assumed by Cicron Cicron intends to pay $225 million in cash and common stock forall f Browns stock in addition to ssuming all of Browns debt. Currently, the market price of Browas common stock is $200 million. Selected items from Browns financial data are as follows 2004 2005 2008 2007 THEREAFTER (MILLIONS Net sales Adninstrative and selling expenses25 25 25 30 30 Depreciation Capital expenditures $260 $265 $280 $20 $300 15 17 18 23 30 18 20 In adition, the cost afgoods sold runs 50percent of sales and the marginal tar rat is 34 percent Computethe net present ahue of the acquisition. Required: a. Compute NPV of the acquisition. b. Make decision to acquire or not. Explain
Answer #1

Estimated Incremental Cash Flows

2004 2005 2006 2007 Thereafter
Net Sales $260 $265 $280 $290 $300
Cost of Goods Sold 130 132.5 140 145 150
Administrative and selling expenses 25 25 25 30 30
Earnings before depreciation and interest 105 107.5 115 115 120
Depreciation 15 17 18 23 30
Earnings before interest and taxes 90 90.5 97 92 90
Taxes 30.6 30.77 32.98 31.28 30.60
Net income 59.40 59.73 64.02 60.72 59.40
+Depreciation 15 17 18 23 30
-Capital expenditures 22 18 18 20 22
Free cash flow 52.40 58.73 64.02   63.72 67.40

Step 1. Estimate the incremental cash flows from the target firm, including the synergistic flows, such as any possible flows from tax credits. This estimation for Brown Corporation is calculated below. Here we are assuming that any cash flows after 2007 will be constant at $67.40 million. Also, we subtract any funds that must be reinvested in the firm in the form of capital expenditures that are required to support the firm’s increasing profits.

Step 2. Determine an appropriate risk-adjusted discount rate for evaluating Brown Corporation . Here we will use the weighted cost of capital (kwacc) for Brown Corporation as our discount rate, where the weighted cost of capital is calculated as kwacc = wdkd (1 − T) + wckc where Wd, Wc = the percentage of funds provided by debt and common, respectively, and Kd , Kc = the cost of debt and common, respectively, and T = the corporate tax rate. For Brown Corporation,

kwacc = (.25)(.08) + (.75)(.22) = .185 or 18.5%

Step 3. Next we must calculate the present value of the incremental cash flows expected from the target firm, using 18.50% as the discounting factor

present value of all cash flows = present value of 2004 to 2007 cash flows + present value of cash flows after 2007,

where the present value of cash flows for 2004 through 2007 would be $156.721 million, determined as follows:

$52.40/(1+0.1850) +$58.73/(1+0.1850)2+ $64.02(1+0.1850)3+$63.72(1+0.1850)4 = $156.721 million

and the present value of the $67.40 million cash flow stream, beginning in 2008, is computed to be $277.921 million.

$67.40/(0.1850)*1/(1.1850)4 = $ 184.493 million

Thus, the present value of the free cash flows associated with the acquisition of Brown Corporation by Cicron Inc. is $341.214 million, that is, the sum of $156.721 million and $184.493 million.

Step 4. Next we estimate the initial outlay associated with the acquisition. As already noted, the initial outlay is defined as the market value of all securities and cash paid out plus the market value of all debt liabilities assumed. In this case, the market value of the assumed debt obligations is $75 million. This amount, along with the acquisition price of $225 million, comprise the initial outlay of $300 million.

Step 5. Finally, the net present value of the acquisition is calculated by subtracting the initial outlay (calculated in step 4) from the present value of the incremental cash flows from the target firm (calculated in step 3):

a) NPVof the acquisition = PV inflows - initial outlay

= $341.214-$300

= $41.214 million

b)Thus, the acquisition should be undertaken because it has a positive net present value.